OpenStudy (anonymous):

solve for y 4y^2-9y+2=0

OpenStudy (anonymous):

y = -2 y = -1/4

OpenStudy (anonymous):

1.2 Factoring 4y2 + 9y + 2 The first term is, 4y2 its coefficient is 4 . The middle term is, +9y its coefficient is 9 . The last term, "the constant", is +2 Step-1 : Multiply the coefficient of the first term by the constant 4 • 2 = 8 Step-2 : Find two factors of 8 whose sum equals the coefficient of the middle term, which is 9 . -8 + -1 = -9 -4 + -2 = -6 -2 + -4 = -6 -1 + -8 = -9 1 + 8 = 9 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, 1 and 8 4y2 + 1y + 8y + 2 Step-4 : Add up the first 2 terms, pulling out like factors : y • (4y+1) Add up the last 2 terms, pulling out common factors : 2 • (4y+1) Step-5 : Add up the four terms of step 4 : (y+2) • (4y+1) Which is the desired factorization

OpenStudy (midhun.madhu1987):

 find y in both the cases.. which are the solutions...